OpenStudy (anonymous):

Divide (x+7) X ----- / --- (2x+14) 10

OpenStudy (anonymous):

Alright, when dividing fractions , there's a certaing rule that you must follow, the question is as follows:\[\frac{x+7}{2x+14} \div \frac{x}{10}\] The second fraction is changed to its reciprocal and the division is switched to multiplication: \[\frac{x+7}{2x+14} \times \frac{10}{x}\] now, multiply the fractions and simplify ^_^ give it a try :)